Landmark Coal operates a mine. During July, the company obtained 500 tons of ore, which yielded 250 pounds of gold and 63,400 pounds of copper. The joint cost related to the operation was $500,000. Gold sells for $325 per ounce and copper sells for $0.85 per pound. Allocate the joint costs using relative weight. With these costs, what is the profit or loss associated with Copper?
Answer #1

1 pound = 16 ounces

Sales value of Gold =Gold sells for $325 per ounce * 250 pounds of gold * 16 ounces

=325 * 250 * 16 = $1,300,000

Sales value of Copper = 63,400 pounds of copper * copper sells for $0.85 per pound

= 63,400 * 0.85 = 53,890

Cost Allocated to copper using relative sales method = Sales value of Copper/(Sales value of Copper+sales value of Gold)* Joint cost related to the operation

=53,890/(53,890+$1,300,000)* $500,000

=19,901.91

Profit associated to Copper=Sales value of Copper - Cost Allocated to copper using relative sales method

=53,890 - 19,901.91= 33,988.09

Thus, Profit associated to copper is $33,988.09